OpenStudy (anonymous):

If the midpoint between (18, y) and (20, -15) is (19, -5), find the value of y.

OpenStudy (anonymous):

can anyone help me with this question..?

OpenStudy (anonymous):

to find mid point we use this lets take mid point(X,Y). and end points as (x1,y1) and (x2,y2)

OpenStudy (anonymous):

and X =(x1+x2)/2 and Y =(y1+y2)/2

OpenStudy (anonymous):

so am i working with (20,-15) and (19,-5) right now?

OpenStudy (anonymous):

end points are (20,-15) and (18 ,y ). here x1= 20 , x2 =18 ,y1= -15 ,y2 = y and X =19 and Y= -5 solve the rest your self

OpenStudy (anonymous):

please someone help me in the question made by me which u can see in the question's list

OpenStudy (anonymous):

math.science if ur not here to help then stop writing in this section

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

but I really need help, I am gonna fail in tomorrow's test

OpenStudy (anonymous):

mid point of x= (x1+x2)/2=(18+20)/2=19 mid point of y= (y1+y2)/2=(y1-15)/2=-5

OpenStudy (anonymous):

yea i got that, but is there more to go?

OpenStudy (anonymous):

sorry i just dont understand how to get y

OpenStudy (anonymous):

y-15=-5*2

OpenStudy (anonymous):

that would be -10

OpenStudy (anonymous):

y=-5*2+15

OpenStudy (anonymous):

this not y you need to take the -15 to the other side

OpenStudy (anonymous):

y=-10+15=5

OpenStudy (anonymous):

so y = 5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok thanks :)